OpenStudy (anonymous):

Which inequality will have a solid boundary line and a shaded region above its graph? x − y ≥ 3 2x − 3y ≤ 3 3y − x < 2 2x + y < 7

OpenStudy (texaschic101):

a solid line means it has an equal sign shaded above the line means it is greater then first we must put these in y = mx + b form x - y >= 3 -y >= -x + 3 y <= x - 3....nope, because this is less then, that means shaded below the line 2x - 3y <= 3 -3y <= -2x + 3 y >= 2/3x - 1....solid line because of equal sign, greater then means shaded above the line....BINGO, this one is your answer. The reason why the others are not even considered is because a solid line requires an equal sign, and the last 2 do not contain an equal sign, therefore, their lines are dashed.

OpenStudy (anonymous):

thank you